# Lennart Berg and Torben Wiedemann

### v1

@@ 6 The parametrisation theorem
Consider the $H_{3}$-graded group $\mathcal{G}\coloneqq(\operatorname{Ch}(D_{6},\mathbb{C}),(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}})$ constructed in 4.10 for $\mathcal{R}\coloneqq\mathbb{C}$ and the faithful parameter system $(A,\mathbb{C}^{2})$ from 6.10. Then there exists a unique $\Delta$-parity map $\eta\colon H_{3}\times\Delta\rightarrow A$ such that $\mathcal{G}$ is parametrised by $(A,\mathbb{C}^{2})$ with respect to $\eta$, the Weyl elements $(w^{H}_{\delta}(1,1))_{\delta\in\Delta}$ from 4.14 and the root isomorphisms $(\tilde{\theta}_{\alpha}^{H})_{\alpha\in H_{3}}$ from 4.10. Further, $\eta$ is complete, braid-invariant and satisfies $\eta(\alpha,\delta)=\eta(-\alpha,\delta)$ for all $\alpha\in H_{3}$ and $\delta\in\Delta$. The values of $\eta$ are given in Figure 5.
@@
Put $w_{\delta}\coloneqq w^{H}_{\delta}(1,1)$ for all $\delta\in\Delta$. For all $\alpha\in H_{3}$ and $\delta\in\Delta$, a straightforward computation (see 2.11) yields a tuple $(\varepsilon,\varepsilon^{\prime})\in\{\pm 1\}^{2}$ such that, for all $x,y\in\mathbb{C}$,
$\tilde{\theta}_{\alpha}^{H}(x,y)^{w_{\delta}}=\tilde{\theta}_{\alpha^{\sigma(\delta)}}^{H}(\varepsilon x,\varepsilon^{\prime}y)\quad\big{(}=\tilde{\theta}_{\alpha^{\sigma(\delta)}}^{H}((\varepsilon,\varepsilon^{\prime}).(x,y))\big{)}.$
We now put $\eta_{\alpha,\delta}\coloneqq(\varepsilon,\varepsilon^{\prime})$. Since $1_{\mathbb{C}}\neq-1_{\mathbb{C}}$, the action of $A$ on $\mathbb{C}^{2}$ is faithful, so the signs $\varepsilon,\varepsilon^{\prime}$ are uniquely determined by $\alpha$ and $\delta$. The results of these computations are given in Figure 5, and we observe that $\eta(\alpha,\delta)=\eta(-\alpha,\delta)$ for all $\alpha\in H_{3}$ and $\delta\in\Delta$.
It remains to verify that the $\Delta$-parity map $\eta:H_{3}\times\Delta\rightarrow A$ has the desired properties. We can read off from Figure 5 that $\eta(\rho_{2},(\rho_{1},\rho_{1}))=(-1,-1)$ and $\eta(\rho_{2},(\rho_{3},\rho_{3}))=(-1,1)$, so $\eta$ is complete by 6.7. Further, braid-invariance holds by 3.15 and the faithfulness of the action of $A$ on $\mathbb{C}^{2}$. (Alternatively, braid-invariance can be derived from Figure 5 with a lot of patience.) ∎
$[0,0,1]$ $(1,1)$ $(-1,-1)$ $(-1,-1)$
$[0,\tau,1]$ $(1,1)$ $(1,-1)$ $(1,1)$
$[0,\tau,\tau]$ $(-1,-1)$ $(1,1)$ $(1,-1)$
$[\tau,\tau,1]$ $(-1,1)$ $(1,1)$ $(-1,-1)$
$[0,1,\tau]$ $(-1,1)$ $(1,-1)$ $(-1,-1)$
$[\tau,\tau,\tau]$ $(1,-1)$ $(-1,-1)$ $(-1,1)$
$[0,1,0]$ $(-1,-1)$ $(-1,-1)$ $(1,-1)$
$[1,1,\tau]$ $(1,-1)$ $(1,1)$ $(-1,1)$
$[\tau,\tau^{2},\tau]$ $(-1,1)$ $(1,1)$ $(-1,-1)$
$[1,1,0]$ $(1,1)$ $(-1,-1)$ $(1,1)$
$[1,\tau^{2},\tau]$ $(-1,-1)$ $(-1,1)$ $(1,-1)$
$[\tau,\tau^{2},\tau^{2}]$ $(1,1)$ $(-1,-1)$ $(1,1)$
$[1,0,0]$ $(-1,-1)$ $(1,1)$ $(1,1)$
$[1,\tau^{2},\tau^{2}]$ $(1,-1)$ $(1,1)$ $(-1,1)$
$[\tau,2\tau,\tau^{2}]$ $(1,1)$ $(-1,1)$ $(1,1)$
Figure 5: Values of the parity map $\eta$ in 6.16, using the notation from 4.6. The values $\eta_{\alpha,\delta}$ for negative roots $\alpha$ are given by the formula $\eta_{-\alpha,\delta}=\eta_{\alpha,\delta}$.
@@
Note that $\eta$ can be defined without reference to $\mathcal{G}$ by declaring that it is given by the values in Figure 5. Further, $\mathcal{G}_{\mathcal{R}}\coloneqq(\operatorname{Ch}(D_{6},\mathcal{R}),(\tilde{U}_{\alpha}^{H})_{\alpha\in H_{3}})$ is parametrised by $(A,\mathcal{R}^{2})$ with respect to $\eta$ for any commutative ring $\mathcal{R}$. However, if $2_{\mathcal{R}}=0_{\mathcal{R}}$, then $(A,\mathcal{R}^{2})$ is not faithful, so there exist multiple parity maps $\eta^{\prime}$ such that $\mathcal{G}_{\mathcal{R}}$ is parametrised by $(A,\mathcal{R}^{2})$ with respect to $\eta^{\prime}$.
Proposition 6.18.
Let $\eta$ be the parity map from 6.16 and let $A\coloneqq\{\pm 1\}^{2}$ act on the root groups $(U_{\alpha})_{\alpha\in H_{3}}$ of $G$ as in 6.12. Then $G$ is square-compatible with respect to $\eta$.
Proof.
In addition to its action on the root groups, we let $A$ act on $\mathbb{C}^{2}$ as in 6.10. Since the group $\mathcal{G}$ from 6.16 is parametrised with respect to $\eta$, we have
$\tilde{\theta}_{\alpha}^{H}(x,y)^{w_{\delta}^{2}}=\tilde{\theta}_{\alpha}^{H}\big{(}\eta(\alpha,(\delta,\delta)).(x,y)\big{)}$
for all $x,y\in\mathbb{C}$. At the same time, the action of $w_{\delta}^{2}$ on $\tilde{\theta}_{\alpha}^{H}(x,y)$ is completely described by 5.20 and 5.21. Since the parameter system $(A,\mathbb{C}^{2})$ is faithful, we conclude: $\eta(\alpha,(\delta,\delta))=(1,1)$ if $\Psi_{\alpha,\delta}$ is of type $A_{1}$ or $A_{1}\times A_{1}$; $\eta(\alpha,(\delta,\delta))=(-1,-1)$ if $\Psi_{\alpha,\delta}$ is of type $A_{2}$; $\eta(\alpha,(\delta,\delta))=(-1,1)$ if $\alpha$ is in involution position with respect to $\delta$ and $\eta(\alpha,(\delta,\delta))=(1,-1)$ if $\alpha$ is in inverted involution position with respect to $\alpha$. Since 5.20 applies to $(G,(U_{\alpha})_{\alpha\in H_{3}})$ as well, the previous statement and the definition of the action of $A$ on $(U_{\alpha})_{\alpha\in H_{3}}$ imply that $G$ is square-compatible with respect to $\eta$. ∎
Conclusion 6.19.
Let $(G,(U_{\alpha})_{\alpha\in H_{3}})$ be an arbitrary $H_{3}$-graded group, let $(w_{\delta})_{\delta\in\Delta}$ be a family of Weyl elements in $G$, let $A\coloneqq\{\pm 1\}^{2}$ denote the twisting group from 6.12 and let $\eta\colon H_{3}\rightarrow A$ denote the parity map from 6.16. Then there exist an abelian group $(\mathcal{S},+)$ on which $A$ acts and a parametrisation $(\theta_{\alpha})_{\alpha\in H_{3}}$ of $G$ by $(A,\mathcal{S})$ with respect to $\eta$ and $(w_{\delta})_{\delta\in\Delta}$. Putting $s^{*}\coloneqq(-1,1).s$ for all $s\in\mathcal{S}$, we can summarise the property of being a parametrisation as follows:
for all $\alpha\in H_{3}$, $\delta\in\Delta$ and $s\in\mathcal{S}$. Note that $(s^{*})^{*}=s$ for all $s\in\mathcal{S}$ because $(-1,1)^{2}=(1,1)$. Further, we define a multiplication on $\mathcal{S}$ by
$a\cdot b\coloneqq\theta_{\rho_{1}+\rho_{2}}^{-1}([\theta_{\rho_{1}}(a),\theta_{\rho_{2}}(b)]).$
It satisfies the distributive laws by 3.8 because $(\rho_{1},\rho_{2})$ is an $A_{2}$-pair. Finally, we define $1_{\mathcal{S}}$ to be the element of $\mathcal{S}$ which satisfies $w_{\rho_{2}}\in U_{-\rho_{2}}\theta_{\rho_{2}}(1_{\mathcal{S}})U_{-\rho_{2}}$, which is uniquely determined by 5.3 (iii). By 5.3 (iv), we have
for all $s\in\mathcal{S}$. A similar computation (see [Wie23, 5.7.12]) shows that $1_{\mathcal{S}}\cdot r=r$ for all $r\in\mathcal{S}$ as well, so that $(\mathcal{S},+,\cdot)$ is a nonassociative ring with unit $1_{\mathcal{S}}$.
Goal 6.21.
We want to show that the multiplication on $\mathcal{S}$ from 6.19 is associative and commutative, that $\mathcal{S}$ decomposes as $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ for a pair of isomorphic ideals $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ of $\mathcal{S}$ and that $(r,s)^{*}=(-r,s)$ for all $(r,s)\in\mathcal{S}_{1}\oplus\mathcal{S}_{2}$. Further, we will show that the commutator relations in $G$ are precisely those given in Figure 4.
@@
All results in this section hold for $H_{4}$-graded groups as well, except that the parity map $\eta$ has to be defined with respect to the $H_{4}$-graded group from 4.23.
@@ 7 The blueprint technique
Notation 7.1.
In this section, Notation 6.1 continues to hold. Further, we define the nonassociative ring $\mathcal{S}$ and the root isomorphisms $(\theta_{\alpha})_{\alpha\in H_{3}}$ as in 6.19.
The blueprint technique was introduced in [Wie23] as a computational tool to prove identities in root graded groups for arbitrary root systems (of rank at least 3). It is inspired by the work of Ronan-Tits on the construction of buildings in [RT87]. In this section, we give a brief summary of the general technique by the example of $H_{3}$-gradings. A more detailed exposition can be found in [Wie23, 6.1]. In Section 8, we will perform the actual blueprint computations for $H_{3}$-graded groups in order to achieve the goal in 6.21.
At first, we have to introduce some notation and terminology.
@@
Let $f$, $g$ be two words over $\Delta$. We say that $f$ is elementary braid-homotopic to $g$ if there exist $\alpha,\beta\in\Delta$ and (possibly empty) words $f_{1},f_{2}$ over $\Delta$ such that $f=f_{1}P_{m}(\alpha,\beta)f_{2}$ and $g=f_{1}P_{m}(\beta,\alpha)f_{2}$ where $m$ denotes the group-theoretic order of $\sigma_{\alpha}\sigma_{\beta}$ in the Weyl group and $P_{m}$ is defined as in 6.5. We say that $f$ and $g$ are braid-homotopic if they lie in the equivalence relation generated by the relation “elementary braid-homotopic”. Further, we say that $f$ is reduced if it does not contain a subword of the form $(\alpha,\alpha)$ for some $\alpha\in\Delta$.
Notation 7.3.
For any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{k})$ over $\Delta$, we put $U_{\bar{\delta}}\coloneqq U_{\delta_{1}}\times\cdots\times U_{\delta_{k}}$.
@@
Lemma 7.6 ([Wie23, 6.2.14]).
Let $\bar{\alpha},\bar{\alpha}^{\prime},\bar{\alpha}^{\prime\prime},\bar{\beta}$ be words over $\Delta$ and let $\varphi$ be a blueprint rewriting rule of type $(\bar{\alpha},\bar{\beta})$. Then $\mathord{\mathrm{id}}\times\varphi\times\mathord{\mathrm{id}}\colon U_{\bar{\alpha}^{\prime}}\times U_{\bar{\alpha}}\times U_{\bar{\alpha}^{\prime\prime}}\rightarrow U_{\bar{\alpha}^{\prime}}\times U_{\bar{\beta}}\times U_{\bar{\alpha}^{\prime\prime}}$ is a blueprint rewriting rule.
The heart of the blueprint technique is the following result. It follows from the fact that the blueprint invariant is injective for any reduced word, which is a consequence of 3.10.
Theorem 7.7 ([Wie23, 6.2.8]).
@@
We assume that for all distinct $i,j\in\{1,2,3\}$, we have a blueprint rewriting rule $\varphi_{ij}$ of type $(P_{m_{ij}}(\rho_{i},\rho_{j}),P_{m_{ji}}(\rho_{j},\rho_{i}))$. For example, we have blueprint rewriting rules $\varphi_{12}\colon U_{121}\rightarrow U_{212}$ and $\varphi_{21}\colon U_{212}\rightarrow U_{121}$. Using these “elementary” blueprint rewriting rules and 7.6, we can define blueprint rewriting rules $\varphi_{i}$ of type $(f_{i},f_{i+1})$ for all $i\in\{1,\ldots,62\}$. Similarly, we can define blueprint rewriting rules $\varphi_{i}^{\prime}$ of type $(f_{i+1},f_{i})$ for all $i\in\{2,\ldots,63\}$.
Now let $y_{1},y_{2},\ldots,y_{15}\in\mathcal{R}$ be arbitrary ring elements and consider the tuple
@@
For all $i\in\{2,\ldots,63\}$, we can successively compute tuples $x_{i}\coloneqq\varphi_{i-1}(x_{i-1})$. Since the identity map on $U_{f_{1}}$ and $\varphi_{62}\circ\cdots\circ\varphi_{1}$ are blueprint rewriting rules, it follows from 7.7 that $x_{1}=x_{63}$. Writing
@@
In practice, we will only compute $x_{2},\ldots,x_{32}$ and, in a separate computation, tuples $x_{63}^{\prime},\ldots,x_{32}^{\prime}$ defined by $x_{63}^{\prime}\coloneqq x_{1}$ and $x_{i}^{\prime}\coloneqq\varphi_{i+1}^{\prime}(x_{i+1}^{\prime})$ for all $i\in\{32,\ldots,62\}$. By the same argument as above, we then have $x_{32}=x_{32}^{\prime}$, which yields 15 identities in $\mathcal{R}$ (which are shorter than the 15 identities in the previous paragraph). We will refer to these identities as the blueprint identities.
@@
The computation in 7.8 will be refered to as the blueprint computation. In order to actually perform this computation, it remains to compute the (elementary) blueprint rewriting rules.
Notation 7.9 (compare 4.20).
For all non-proportional roots $\alpha,\gamma\in H_{3}$ and all $\beta\in\mathord{]\alpha,\gamma[}$, we define a map $\psi_{\alpha,\gamma}^{\beta}\colon\mathcal{R}^{2}\rightarrow\mathcal{R}$ by the property that
for all $x_{1},x_{2},y_{1},y_{2}\in\mathcal{R}$. If $\mathord{]\alpha,\gamma[}=\{\beta\}$, we will write $\psi_{\alpha,\gamma}$ in place of $\psi_{\alpha,\gamma}^{\beta}$.
Notation 7.10.
From now on, we put $\alpha\coloneqq\rho_{2}$, $\varepsilon\coloneqq\rho_{3}$ and we denote the $H_{2}$-quintuple corresponding to the $H_{2}$-pair $(\alpha,\varepsilon)$ by $(\alpha,\beta,\gamma,\delta,\varepsilon)$.
@@
The following computation shows how to arrive at the formulas in 7.11.
Lemma 7.12 ([Wie23, 6.3.4]).
@@
and put $f\coloneqq(\rho_{1},\rho_{2},\rho_{1})$. Then, by 7.4, Figure 5 and (6.20), we have
@@
Since $A=c$, $B=-b-ca$ and $C=a$ by 7.11, we infer that $\varphi_{12}$ is a blueprint rewriting rule. By a similar calculation, the same holds for $\varphi_{13}$. Lastly, note that $\varphi_{21}$ and $\varphi_{31}$ are the inverses of $\varphi_{12}$ and $\varphi_{13}$, respectively. Hence they are blueprint rewriting rules by 7.5. ∎
@@
The proof proceeds just as the one of 7.12. We omit the details. ∎
8 The blueprint computation for $H_{3}$-graded groups
Notation 8.1.
The notation introduced in 7.1 and 7.10 continues to hold. That is, we fix an $H_{3}$-graded group $(G,(U_{\alpha})_{\alpha\in H_{3}})$ and a choice of Weyl elements $(w_{\delta})_{\delta\in\Delta}$ for a root base $\Delta=(\rho_{1},\rho_{2},\rho_{3})$ of $H_{3}$ in standard order. We define $\mathcal{S}$ and $(\theta_{\alpha})_{\alpha\in H_{3}}$ as in 6.19 and maps $\psi_{\alpha,\gamma}^{\beta}$ as in 7.9, and we denote by $(\alpha,\beta,\gamma,\delta,\epsilon)$ the $H_{2}$-quintuple with $\alpha=\rho_{2}$ and $\epsilon=\rho_{3}$.
Using the blueprint rewriting rules from 7.11 and the homotopy cycle from Figure 6, we can perform the blueprint computation for $(G,(U_{\alpha})_{\alpha\in H_{3}})$, following the algorithm in 7.8. This yields 15 blueprint identities involving indeterminates $y_{1},\ldots,y_{15}\in\mathcal{S}$. Using these identities, we will show in this section that $\mathcal{S}$ is an associative commutative ring (see 8.3, 8.26) which decomposes as a direct sum $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ of two isomorphic ideals $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ (see 8.21, 8.24), and we will also compute explicit formulas for the commutation maps $\psi_{\xi,\zeta}^{\rho}$ from 6.19.

### v2

@@ 7 Standard coordinatisations and the Steinberg presentation
$(\operatorname{St}_{H_{\ell}}(\mathcal{R}),(U_{\alpha}^{\text{St}})_{\alpha\in H_{\ell}})$ is an $H_{\ell}$-graded group with standard coordinatisation $(\theta_{\alpha}^{\text{St}})_{\alpha\in H_{\ell}}$.
@@
By 7.3, 7.6 and 7.7, $(U_{\alpha}^{\text{St}})_{\alpha\in H_{\ell}}$ is an $H_{\ell}$-grading of $\operatorname{St}_{H_{\ell}}(\mathcal{R})$ with root isomorphisms $(\theta_{\alpha}^{\text{St}})_{\alpha\in H_{\ell}}$. By similar arguments as in 7.7 (that is, projecting into the appropriate Chevalley group and using that the projection is injective on positive systems), the action of the Weyl elements on the root groups in $\operatorname{St}_{H_{\ell}}(\mathcal{R})$ satisfies the same formulas as the corresponding Weyl elements in Chevalley groups. Hence Axiom 6.28 ii is satisfied. Thus $(\theta_{\alpha}^{\text{St}})_{\alpha\in H_{\ell}}$ is a standard coordinatisation. ∎
Proposition 7.9.
Let $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ be an $H_{\ell}$-graded group with a standard coordinatisation $(\theta_{\beta})_{\beta\in H_{\ell}}$ by $\mathcal{R}\times\mathcal{R}$ and let $\alpha\in H_{\ell}$. Then the map
×××$\mathcal{R}^{\times}\times\mathcal{R}^{\times}\rightarrow M_{\alpha},(r,s)\mapsto w_{\alpha}(r,s)\coloneqq\theta_{-\alpha}(-r^{-1},-s^{-1})\theta_{\alpha}(r,s)\theta_{-\alpha}(-r^{-1},-s^{-1})$
is a bijection. Here $\mathcal{R}^{\times}$ and $M_{\alpha}$ denote the sets of invertible elements in $R$ and of $\alpha$-Weyl elements in $G$, respectively.
Proof.
A similar assertion is proven in [Wie24, 5.6.6] for $\Phi$-graded groups where $\Phi$ is of type $A$, $D$ or $E$, and the same kinds of arguments apply in our situation. For the reader’s convenience, we repeat the main ideas. Note that for all $r,s\in\mathcal{R}^{\times}$, $w_{\alpha}^{\text{St}}(r,s)$ is a Weyl element in $\operatorname{St}_{H_{\ell}}(\mathcal{R})$ by 7.7 and hence its image $w_{\alpha}(r,s)$ in $G$ is a Weyl element in $G$.
At first, let $\alpha\in H_{\ell}$ and let $w_{\alpha}$ be any $\alpha$-Weyl element in $G$. By 5.3 (iii) there exist unique $a,b,c,d\in\mathcal{R}$ such that
$w_{\alpha}=\theta_{-\alpha}(a,b)\theta_{\alpha}(c,d)\theta_{-\alpha}(a,b).$
Choose a root $\gamma\in H_{\ell}$ such that $(\alpha,\gamma)$ is an $A_{2}$-pair. Then by the commutator relations in $\operatorname{St}_{H_{\ell}}(\mathcal{R})$, there exist signs $\varepsilon_{1},\varepsilon_{2}\in\{\pm 1\}$ such that
$[\theta_{\gamma}(r_{1},r_{2}),\theta_{\alpha}(s_{1},s_{2})]=\theta_{\alpha+\gamma}(\varepsilon_{1}r_{1}s_{1},\varepsilon_{2}r_{2}s_{2})$
for all $r_{1},r_{2},s_{1},s_{2}\in\mathcal{R}$. By an application of 5.3 (iv), we infer that
$\theta_{\gamma}(r,s)^{w_{\alpha}}=[\theta_{\gamma}(r,s),\theta_{\alpha}(c,d)]=\theta_{\alpha+\gamma}(\varepsilon_{1}rc,\varepsilon_{2}sd).$
for all $r,s\in\mathcal{R}$. Since the map ×$\mathcal{R}\times\mathcal{R}\rightarrow U_{\alpha+\gamma},(r,s)\mapsto\theta_{\gamma}(r,s)^{w_{\alpha}}$ is bijective, it follows that $c$ and $d$ are invertible. Similar arguments applied to the formula
$\theta_{\gamma}(1_{\mathcal{R}},1_{\mathcal{R}})=\big{[}[\theta_{\gamma}(1_{\mathcal{R}},1_{\mathcal{R}}),\theta_{\alpha}(c,d)],\theta_{-\alpha}(a,b)]^{-1}$
from 5.3 (iv) show that $c^{-1}=-a$ and $d^{-1}=-b$. (The precise computation, which we skip, relies on the knowledge of the signs $\varepsilon_{1}$, $\varepsilon_{2}$.) This shows that the map in the assertion is surjective. By 5.3 (iii), it is also injective. ∎
@@
Let $\Phi$ be any root system. An RGD-system of type $\Phi$ is (essentially) a $\Phi$-graded group $(G,(U_{\alpha})_{\alpha\in\Phi})$ such that $U_{\alpha}^{\sharp}=U_{\alpha}\setminus\{1_{G}\}$. These objects were introduced by Tits in [Tit92], and they are (essentially) in bijective correspondence with thick Moufang buildings of type $\Phi$ (see [AB08, Section 7.8] for details). It follows from 7.9 that an $H_{\ell}$-graded group with a standard coordinatisation cannot be an RGD-system (because $\mathcal{R}\times\mathcal{R}$ contains non-invertible elements, such as $(x,0)$). Since every $H_{\ell}$-grading has a standard coordinatisation by our main result 9.33, it follows that there exist no RGD-systems (and hence no thick Moufang buildings) of types $H_{3}$ and $H_{4}$. This is a classical result of Tits [Tit77, 3.5], which he proved by showing the analogous result for type $H_{2}$ [Tit76, Theorème 1].
@@
Let $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$ be an $H_{\ell}$-graded group with a standard coordinatisation $(\theta_{\alpha})_{\alpha\in H_{\ell}}$ by $\mathcal{R}\times\mathcal{R}$. It follows from 7.5 that for any positive system $\Pi$ in $H_{\ell}$, the subgroup $U_{\Pi}$ of $G$ is isomorphic to the subgroup $\tilde{U}_{\Pi}^{H}$ of $\operatorname{Ch}(X,\mathcal{R})$. Thus $G$ cannot be “substantially different” from $\operatorname{Ch}(X,\mathcal{R})$, but it need not be isomorphic to $\operatorname{Ch}(X,\mathcal{R})$. For example, $G$ could be the adjoint Chevalley group of type $D_{6}$ as well as the simply-connected Chevalley group of type $D_{6}$, which are not isomorphic. We do not prove that $G$ is (isomorphic to) a Chevalley group, but we are also not aware of an example of an $H_{\ell}$-graded group which is provably not a Chevalley group.
@@ 8 The blueprint technique
In [RT87], Ronan-Tits introduced the notion of blueprints as a tool for constructing (and hence proving the existence of) buildings of certain types. More precisely, they show that if certain computations involving rewriting rules and a homotopy cycle of the longest element in the Weyl group lead to a certain result, then a blueprint can be used to construct a building. What we refer to as the “blueprint technique” is the reversal of this idea: Starting from a root graded group $G$ (which should be thought of as a group acting on a “generalised building”) with a parametrisation (in the sense of 6.11), one can perform the same kind of computations to obtain identities that are satisfied by the commutation maps (defined as in 3.17) in $G$. This application of the ideas in [RT87] was first described in [Wie24].
In this section, we give a brief summary of the general technique by the example of $H_{\ell}$-gradings. A more detailed exposition can be found in [Wie24, 6.1]. In Section 9, we will perform the actual blueprint computations for $H_{\ell}$-graded groups in order to achieve the goal in 6.27.
We begin by introducing some notation and terminology.
Notation 8.1.
In this section, we fix $\ell\in\{3,4\}$ and an $H_{\ell}$-graded group $(G,(U_{\alpha})_{\alpha\in H_{\ell}})$. We also choose a root base $\Delta$ of $H_{\ell}$ in standard order, which we denote by $(\rho_{0},\rho_{1},\rho_{2},\rho_{3})$ if $\ell=4$ and by $(\rho_{1},\rho_{2},\rho_{3})$ if $\ell=3$. We will always consider $H_{3}$ as the subsystem $\Psi_{\rho_{1},\rho_{2},\rho_{3}}$ of $H_{\ell}$. Finally, we fix a family $(w_{\delta})_{\delta\in\Delta}$ of Weyl elements and define the nonassociative ring $\mathcal{S}$ and the root isomorphisms $(\theta_{\alpha})_{\alpha\in H_{3}}$ as in 6.24.
Remark 8.2.
Nearly all computations in this section and the following one take place in the subgroup of $G$ generated by $(U_{\alpha})_{\alpha\in H_{3}}$. Only in 9.30 and 9.32 will we have to leave the $H_{3}$-subsystem to ensure that the commutator relation for $(\rho_{0},\rho_{1})$ and the Weyl element $w_{\rho_{0}}$ behave as they should.
@@
A word $f$ over $\Delta$ is reduced if it is not homotopic (in the sense of 6.17) to shorter word.
Notation 8.4.
For any word $\bar{\delta}=(\delta_{1},\ldots,\delta_{k})$ over $\Delta$, we put
$U_{\bar{\delta}}\coloneqq U_{\delta_{1}}\times\cdots\times U_{\delta_{k}}.$
@@
Lemma 8.7 ([Wie24, 6.2.14]).
Let $\bar{\alpha},\bar{\alpha}^{\prime},\bar{\alpha}^{\prime\prime},\bar{\beta}$ be words over $\Delta$ and let $\varphi$ be a blueprint rewriting rule of type $(\bar{\alpha},\bar{\beta})$. Then $\operatorname{id}\times\varphi\times\operatorname{id}\colon U_{\bar{\alpha}^{\prime}}\times U_{\bar{\alpha}}\times U_{\bar{\alpha}^{\prime\prime}}\rightarrow U_{\bar{\alpha}^{\prime}}\times U_{\bar{\beta}}\times U_{\bar{\alpha}^{\prime\prime}}$ is a blueprint rewriting rule.
The heart of the blueprint technique is the following result. It follows from the fact that the blueprint invariant is injective for any reduced word, which is a consequence of 3.11.
Theorem 8.8 ([Wie24, 6.2.8]).
@@
We assume that for all distinct $i,j\in\{1,2,3\}$, we have a blueprint rewriting rule $\varphi_{ij}$ of type $(P_{m_{ij}}(\rho_{i},\rho_{j}),P_{m_{ji}}(\rho_{j},\rho_{i}))$. For example, we have blueprint rewriting rules $\varphi_{12}\colon U_{121}\rightarrow U_{212}$ and $\varphi_{21}\colon U_{212}\rightarrow U_{121}$. Using these “elementary” blueprint rewriting rules and 8.7, we can define blueprint rewriting rules $\varphi_{i}$ of type $(f_{i},f_{i+1})$ for all $i\in\{1,\ldots,62\}$. Similarly, we can define blueprint rewriting rules $\varphi_{i}^{\prime}$ of type $(f_{i+1},f_{i})$ for all $i\in\{2,\ldots,63\}$.
Now let $y_{1},y_{2},\ldots,y_{15}\in\mathcal{S}$ be arbitrary ring elements and consider the tuple
@@
For all $i\in\{2,\ldots,63\}$, we can successively compute tuples $x_{i}\coloneqq\varphi_{i-1}(x_{i-1})$. Since the identity map on $U_{f_{1}}$ and $\varphi_{62}\circ\cdots\circ\varphi_{1}$ are blueprint rewriting rules, it follows from 8.8 that $x_{1}=x_{63}$. Writing
@@
In practice, we will only compute $x_{2},\ldots,x_{32}$ and, in a separate computation, tuples $x_{63}^{\prime},\ldots,x_{32}^{\prime}$ defined by $x_{63}^{\prime}\coloneqq x_{1}$ and $x_{i}^{\prime}\coloneqq\varphi_{i+1}^{\prime}(x_{i+1}^{\prime})$ for all $i\in\{32,\ldots,62\}$. By the same argument as above, we then have $x_{32}=x_{32}^{\prime}$, which yields 15 identities in $\mathcal{S}$ (which are shorter than the 15 identities in the previous paragraph). We will refer to these identities as the blueprint identities.
@@
The computation in 8.9 will be refered to as the blueprint computation. In order to actually perform this computation, it remains to compute the (elementary) blueprint rewriting rules.
Notation 8.10.
From now on, we put $\alpha\coloneqq\rho_{2}$, $\varepsilon\coloneqq\rho_{3}$ and we denote the $H_{2}$-quintuple corresponding to the $H_{2}$-pair $(\alpha,\varepsilon)$ by $(\alpha,\beta,\gamma,\delta,\varepsilon)$. We will also make frequent use of Notation 3.17 for commutation maps (with respect to $(\theta_{\zeta})_{\zeta\in H_{\ell}}$).
@@
The following computation shows how to arrive at the formulas in 8.11.
Lemma 8.12 ([Wie24, 6.3.4]).
@@
and put $f\coloneqq(\rho_{1},\rho_{2},\rho_{1})$. Then, by 8.5, Figure 4 and (6.25), we have
@@
Since $A=c$, $B=-b-ca$ and $C=a$ by 8.11, we infer that $\varphi_{12}$ is a blueprint rewriting rule. By a similar calculation, the same holds for $\varphi_{13}$. Lastly, note that $\varphi_{21}$ and $\varphi_{31}$ are the inverses of $\varphi_{12}$ and $\varphi_{13}$, respectively. Hence they are blueprint rewriting rules by 8.6. ∎
@@
The proof proceeds just as the one of 8.12. We omit the details. ∎
9 The blueprint computation for $H_{3}$-graded
groups
Notation 9.1.
The notation introduced in 8.1 and 8.10 continues to hold.
Using the blueprint rewriting rules from 8.11 and the homotopy cycle from Figure 6, we can perform the blueprint computation for $(G,(U_{\alpha})_{\alpha\in H_{3}})$, following the algorithm in 8.9. This yields 15 blueprint identities involving indeterminates $y_{1},\ldots,y_{15}\in\mathcal{S}$. Using these identities, we will show in this section that $\mathcal{S}$ is an associative commutative ring (see 9.3, 9.26) which decomposes as a direct sum $\mathcal{S}=\mathcal{S}_{1}\oplus\mathcal{S}_{2}$ of two isomorphic ideals $\mathcal{S}_{1}$, $\mathcal{S}_{2}$ (see 9.21, 9.24), and we will also compute explicit formulas for the commutation maps $\psi_{\xi,\zeta}^{\rho}$ from 6.24.